OpenStudy (mikezack123):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer.

OpenStudy (mikezack123):

Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (mikezack123):

@amistre64 @hartnn @whpalmer4 @Preetha

OpenStudy (mikezack123):

@Luigi0210 @magbak @experimentX @cwrw238

OpenStudy (mikezack123):

@texaschic101 @phi @MathsPro @beccaboo333 @AccessDenied

OpenStudy (beccaboo333):

I wish I could help, but I am not good with math.

OpenStudy (mikezack123):

@beccaboo333 it okay, thanks anyways :)

OpenStudy (amistre64):

whats your explanation?

OpenStudy (mikezack123):

what do you mean? @amistre64

OpenStudy (amistre64):

did you read the question?

OpenStudy (mikezack123):

yes why?

OpenStudy (amistre64):

if you cant determine part1 yet, then how would you work part 2?

OpenStudy (mikezack123):

i didn't do any part of the question yet

OpenStudy (amistre64):

im trying to get you to develop a solution without actually giving you the answers. guidance as opposed to direct answers

OpenStudy (amistre64):

try to work the question then ....

OpenStudy (mikezack123):

i don't know where to start thats why i posted it on this website

OpenStudy (anonymous):

i wish i could help but i can't figure it . sorry.

OpenStudy (amistre64):

id start with part 2 ... just becuse i know how i would simplify the problem

OpenStudy (mikezack123):

@princess6324 thanks anyways :)

OpenStudy (amistre64):

i think the 2x part is right, since the sqrts can all simplify to a 2x part and factor out a complete square

OpenStudy (amistre64):

3sqrt(2x) + xsqrt(8x) - 5sqrt(18x) and got -10x sqrt(2x) putting it in the format migh thelp me see it better ... flipping between windows and trying to keep things staight is hard on these old eyes

OpenStudy (amistre64):

can you factor out the perfect squares from out of the sqrts?

OpenStudy (anonymous):

You welcome. Next time i am sure i am going to help you?

OpenStudy (whpalmer4):

"Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. " \[3\sqrt{2x} + x\sqrt{8x} - 5\sqrt{18x} = -10x\sqrt{2x}\] Can you factor all of the things under the square root signs? Are there any perfect squares which can be factored out?